OpenStudy (anonymous):

can i just get answer please Write the equation of a line that goes through the points (0, 2) and (-3, 2)

OpenStudy (anonymous):

Nope! ^_^ use the point slope formula to first find the slope then plug in the slope into point slope formula along with a pair of (x,y) coordinates

OpenStudy (anonymous):

Point slope formul is: y2 - y1/ x2 - x1

OpenStudy (anonymous):

equal to m and m being the slope

OpenStudy (anonymous):

all you do is plug in the numbers! Give it a try, it'll save so much more time than typing each question out and asking for answer ;)

OpenStudy (anonymous):

you plug into y=mx+b?

OpenStudy (anonymous):

y=x-3

OpenStudy (anonymous):

@Weech31912

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

nope just y=x-3

OpenStudy (anonymous):

thats the answer or what?

hartnn (hartnn):

y= x-3 is incorrect.

OpenStudy (anonymous):

oh, well thats the answer i got

hartnn (hartnn):

@Weech31912 first find slope(m) as @Yellowpanda said, what you get m=.... ?

OpenStudy (anonymous):

0

hartnn (hartnn):

thats correct, now use the point slope form, y-y1 = m (x-x1) x1 = 0, y1= 2 , m=0 plug in values, what u get ?

OpenStudy (anonymous):

y-2=0(x-0)

hartnn (hartnn):

yes, that would be y-2 =0 right ? so the equation of your line is just y=2 thats it!

OpenStudy (anonymous):

oh i c i was missin that formula